OpenStudy (anonymous):

Determine the circumference of a circle with a radius of 6.8. Round your answer to the tenths place.

OpenStudy (solomonzelman):

\[C=2πr~~~~~OR~~~~~C=dπ\] C - Circumference π - is the pi, or approximately 3.14 r - Radius d - Deameter

OpenStudy (solomonzelman):

\[π=3.14~~~~~~~~~~~\And~~~~~~~~~~~~r=6.8~~,~~~~~~~~~~~~~Plug~~this~~i n t o:~~~C=2 \times π \times r\]

OpenStudy (anonymous):

Thanks! :) I'll try it.

OpenStudy (solomonzelman):

tell me what you get when you're done. (nearest 10th is a first decimal place for example, \[32.\color{red} {4}267....\]

OpenStudy (anonymous):

I did 2*3.14*6.8 and it =42.7

OpenStudy (solomonzelman):

YES! CORRECT!

OpenStudy (anonymous):

yayyyy!! thanks so muchhh!

OpenStudy (solomonzelman):

Anytime!